Suppose a researcher is interested in predicting anxiety level based on the amount of caffeinated beverages consumed a day. Data collected from 5 college students is shown below. Be sure to labe your answers with the appropriate letter and show all your work (leach mathematical step) for each answer Caffeinated Anxiety M Beverages (X) 10 Calculate the covariance between the two variables Clauatethe siandrd deiatono ech arable andthe corelaton betwen thevariales llt hesoge as i onpleterersin reii equation. d tnewperson drils afeiated verages wha ould you prdi hi aey lee to be HINT Yuneed tous the regression quationto answer this question.
Answer #1

S.No 1 2 2 1 0 4. EX 10.0000 9.0000 4.0000 1.0000 0.0000 16.0000 81.0000 64.0000 36.0000 64.0000 100.0000 27.00 16.00 6.00 0.00 40.00 10 total 41.0000 30.0000 345.0000 89.0000 5.0000 2.0000 8.2000 1.5811 1.4832 1.7500 0.7462 s,-(V(Zx-(Zx)2/n)/(n-1)) Covxys(ΣΧΥ_ (ΣΥΣΥ)/n)/(n-1)- from above slope- 0.700 intercept- 6.800 hence line equation: 6.800 0.700 x

a)

Covariance =1.75

b)

std deviation of x =1.5811

std deviation of y =1.4832

correlation coefficient r=0.7462

c)slope=0.70

intercept =6.80

line equation: Yhat=6.80+0.70*x

d)

predicted value =6.80+0.7*6=11